OpenStudy (anonymous):

A skier, traveling down a hill has a uniform acceleration of +1.3 m/s^2. A)if he starts from rest at the top of the hill and travels 50 m to the bottom, how long(how much time) does it take? B)How far does he travel in that time? i'm just of which equations to use.

OpenStudy (anonymous):

Are you just looking for what equations to use?

OpenStudy (anonymous):

u can use s= ut + 1/2 at^2...........u=0 a=1.3 and s= 50..fint t abt bpart...i dint really get it...the distance travelled is already given

OpenStudy (anonymous):

It might be more involving that. Skier is moving along a slope, he is moved only by gravitational potential energy. By the given acceleration you can get the slope angle: \[\sin \theta = acceleration G / accelerationSLOPE\] You can solve for the angle so you can use the equation or substitute the preceding equation in it \[X(t) = V(0) t \sin \theta+ (a t^2 \sin \theta)/2\] This will give you how far on the slope. Time you'll get by solving for (t) in \[X(50) = V(0) t + (a t^2 ) /2\] Remember V(0) is 0. If non of the above is certain, then the professor should give more time to paraphrase the problems in a more unambiguous way.